Instead of spending the $50 a month you receive, you decided to invest it in your money market account at your bank. Assuming you will earn an average of 5% each year and inflation is forecasted to 2.5% a year, how much have you effectively earned after ten years? (Hint: must use a two formulas to solve this question) In other words, after you have factored out the effects of inflation, how much purchasing power do you have?

Show all work.
Answer #1

we know that the investment is 50$ monthly

Time period of 10 years

yearly average earnings is 5%

number of payments = no. of months in year * no. of years = 12 * 10 = 120

interest rate = 5/12 = 0.416

Then total outcome after 10 years = $7,764

effect of inflation at rate 2.5 = (7764) / (1 + .025 ) 10 = 6065

Then the purchasing power is that you will buy 6065$ good will be bought today with 7764$ after 10 years

HOPE THIS HELPS
PLEASE COMMENT